you deposit $1600 in a bank account. Find the balance after 3 years for each of the following situations:
a) the account pays 2.5% annual interest compounded monthly
b) the account pays 1.75% annual interest compounded quarterly
c) the account pays 4% annual interest compounded yearly

Answer by LinnW(1048)

A = P(1 + r/period )^(period*years)
A is amount in the account after years with compounding at period
a) A = 1600*(1 + 0.025/12)^(12*3) Input into excel or calculator
1724.48
b) A = 1600*(1 + 0.0175/4)^(4*3)
1686.05
c) A = 1600*(1 + 0.04/1)^(1*3)
1799.78